Please help solve this: -(11-2x)=5(x+8)
Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

distributing __ -11+2x = 5x+40

subtracting 2x+40 __ -51 = 3x

dividing by 3 __ -17 = x

Answer by Alan3354(69443)

Please help solve this: -(11-2x)=5(x+8)
--------------
-(11-2x)=5(x+8)
-11+2x = 5x + 40
2x = 5x + 51
-3x = 51
x = -17